A circle passes through the origin at point (3,0) and (0,4). What is the radius of the circle?
Answer by checkley77(12844)

IF I READ THE PROBLEM CORRECTLY. YOU HAVE A RIGHT TRIANGLE WITH THREE POINTS (0,0), (3,0) & (0,4)
ONE SIDE IS 3, THE OTHER SIDE IS 4.
THUS THE LENGTH OF THE HYPOTENUSE WOULD BE THE RADIUS OF THIS CIRCLE.
3^2=4^2=R^2
9+16=C^2
25=C^2
C=5 ANSWER FOR THE RADIUS.